OpenStudy (anonymous):

***Please help me i really need it**** Sandy works at a clothing store. She makes $9 per hour plus earns 15% commission on her sales. She worked 71 hours over the last two weeks and had a total of $2,690 in sales before taxes. Which of the following is closest to how much she will earn in hourly wages and commission for those two weeks? A.499 B.639 C.404 D.1,043

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

help please!

OpenStudy (paki):

15% commission on 9$ = 1.35$ per hour= 1.35*24= 32.4$ per 24 hours= 32.4*14= 453.6 for two weeks.... and 453.6+9= 462$... which is nearly of 499$ ....